Question:
You can bike to your aunt's house at 15mph. Your bike has a flat and you have to walk back at 6mph. If your total traveling time was 2 hours, how long did it take you to ride your bike to your aunt's house?
Distance - rate x time
I think the correct equation is:
15x = 6 (2-x)
15x = 12 - 6x
21x = 12
x = 12/21 or 4/7
Is that 4/7 of an hour? 34 minutes and 29 seconds
So, is the answer "It takes 34 minutes to ride to aunt's house"??

Answer by richwmiller(17219)

x+y=2
x=2-y
15x=6y
15(2-y)=6y
30-15y=6y
30=21y
10=7y
y=10/7= 1 3/7
x=14/7-10/7=4/7 hr 34 min 17 sec